Calculating Average Operating Assets, Margin, Turnover, Return on Investment (ROI) Forchen, Inc., provided the following info. What if operating income for the Small Appliances Division was $2,496,240? How would that affect average operating assets?
Answer #1
A. Average Operating Assets:
Average operating Assets = Beginning Assets + Ending Assets / 2
Small Appliances Division = 6,394,000 + 7,474,000 / 2 $ 6,934,000
Cleaning Products Division = 5,600,000 + 6,000,000 / 2 $ 5,800,000
B. Margin :
Margin = Operating Income/ Sales
Small Appliances Division (3,744,360 / 41,604,000) 9%
Cleaning Products Division (1,160,000 / 29,000,000) 4%
Small Appliances Division (41,604,000 / 6,934,000) 9%
Cleaning Products Division (29,000,000 / 5,800,000) 5.18
D . Return On Investment :
ROI = Operating Income / Average Operating Assets
Small Appliances Division (3,744,360 / 6,934,000) 54%
Cleaning Products Division (1,160,000 / 5,800,000 ) 20%
Margin = Net Income / Sales = (2496240/41604000) 6.00%
Margin lowered by 2.00%
ROI = Net Income / Average Assets = (2496240/6934000) 36.00%
ROI lowered by 18.00%